morganislame:

What is the slope of the line whose equation is 2x−4y=10 ?

Angle:

Hello @morganislame an easy way to do this question is to consider the "slope intercept form" which is y = mx + b where m is the slope

Angle:

so what we want to do is try to get y alone to one side

morganislame:

okay, gotcha

Angle:

do you know how to do that? I can guide you through it if you'd like :)

morganislame:

That would be great. This is a new concept for me, so its a little bit confusing. Thank you!

Angle:

If we want to get "y" alone, we have to add/subtract/multiply/divide stuff to detach the "stuff" crowding the "y" 2x - 4y = 10 ^ we have 2x crowding the y so we try to subtract 2x first 2x - 4y = 10 -2x - 2x


-4y = 10 - 2x do you see how I got that? :)

morganislame:

Yes, i do. That makes sense :)

Angle:

Great! Then -4y = 10 - 2x ^^ we have this "-4" crowding the y but it is attached with multiplication, so we should use division to separate it divide both sides by -4 (-4y)/(-4) = (10 - 2x)/(-4) becomes y = (10 - 2x)/(-4) \[y = \frac{ 10 - 2x }{ -4 }\] can you try simplifying that? :)

Angle:

hint: \[y = \frac{10 - 2x}{-4} = \frac{10}{-4} + \frac{-2x}{-4}\]

morganislame:

I understand it now, and i got the right answer! Thanks for giving me that quick review. :)

Angle:

I'm glad I could help @morganislame Feel free to tag me with @Angle if you need any more help!